x-5y = 2
11x+13y=22

Answer by gems(12)

x - 5y = 2
by substitution, x = 2 + 5y
Use in 2nd equation 11x + 13y = 22
11 (2 + 5y) + 13y = 22
22 + 55y + 13y = 22
68y = 22 - 22
68y = 0
y = 0
If y = 0
x - 5y = 2
x - 5(0) = 2
x - 0 = 2
x = 2
Therefore x = 2, and y = 0